OpenStudy (anonymous):

What is the solution to the system of equations? 6x – 9y = 16 2x – 3y = 7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

me can multiply the bottom by -3 and get 6x – 9y = 16 -6x +9y = -21 -------------- this shows us that these are parallel lines with nothing in common

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

u have to put a line through this= then u can get it